# Distribution Grids May Be a Barrier to Residential Electrification

### v2

@@ RESULTS
Fig. 5 shows heat maps of the estimated reductions in grid reinforcement requirements in each county under the four demand-side management strategies. While heat-pump water heaters use much less electricity than resistance water heaters, the top left heat map in Fig. 5 shows that they have little impact on grid reinforcement requirements due to the relatively small contribution of water heating to peak demand. By contrast, the heat maps with less thermal demand from buildings (top right) and with ground-source heat pumps (bottom right) show significant reductions in grid reinforcement requirements, particularly in the coldest regions. The heat map with device coordination (bottom left) shows modest reductions in grid reinforcement requirements. However, these reductions are more evenly distributed across the country because, for example, coordinating electric vehicle charging can reduce peak demand anywhere, regardless of climate.
Fig. 6 shows nationwide grid reinforcement cost reductions from the four individual strategies and from combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$522 billion without demand-side management. Individually (top set of bars), the four strategies reduce costs by 4%, 13%, 36%, and 43%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication, rather than equipment upgrades – reduces costs more than switching to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of heat-pump water heaters (4% individually), less thermal demand from buildings (13%), and ground-source heat pumps (36%) reduces costs by 46%. The combination of all four demand-side management strategies reduces nationwide grid reinforcement costs by 77%, to a total of \$119 billion (\$950 per household).
@@ DISCUSSION
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can increase peak aggregate demand by a factor of five or more in the coldest counties. Without demand-side management, distribution grid reinforcement could cost \$320 to \$720 billion nationwide, or \$2,600 to \$5,800 per household, but smart electrification could eliminate up to three-quarters of these costs.
@@
This section discusses the sensitivity of nationwide distribution grid reinforcement requirements to several modeling assumptions, including heat pump efficiencies, whether households lower their heating temperature setpoints overnight, heat pump sizing protocols, and the choice of weather data. The Supplementary Information document discusses the sensitivity analyses in more detail.
@@
Heat pump efficiencies significantly influence grid reinforcement requirements. In cold weather, heat pumps become less efficient and may require electric resistance backup to meet demand. The central cases presented above assumed that heat pumps in cold regions met the United States Department of Energy’s Cold-Climate Heat Pump Challenge specifications 34. However, in a sensitivity study where heat pumps in cold regions have efficiencies that are similar to today’s equipment, nationwide distribution grid reinforcement requirements increase by about 18%.
@@
Lowering space heating temperature setpoints overnight can increase grid reinforcement requirements. While lowering setpoints can save energy overnight, warming the thermal mass of a building back up in the morning can require high power and potentially trigger the use of resistance backup heat. In the central cases presented above, users did not alter setpoints overnight. However, in a sensitivity study where users lower setpoints overnight, with randomized but typical setpoint adjustment magnitudes and timings, nationwide distribution grid reinforcement requirements increase by about 18%.
@@
Heat pump sizing has a comparatively small influence on grid reinforcement requirements. In the central cases presented above, we sized heat pumps to meet the larger of the design cooling load and the design heating load, capping the size at the maximum capacity currently available on the market. If heating demand in very cold weather exceeded the heat pump’s capacity, electric resistance backup heat turned on to maintain space temperatures. An alternative approach would be to size heat pumps in colder climates to meet only the design cooling load, relying much more heavily on backup heat in cold weather. We simulated this alternative approach in a sensitivity study and found that it increases nationwide grid reinforcement requirements by about 8%. In the coldest regions, however, sizing heat pumps to the design cooling loads increases grid reinforcement requirements by about 12%.
@@
The sensitivities of grid reinforcement requirements to modeling choices on heat pump efficiency (about 18%), overnight temperature setpoint adjustments (about 18%), and heat pump sizing (about 8%) are more than additive. In a sensitivity study including all three of the alternative modeling choices discussed above, nationwide grid reinforcement requirements increase by about 50%.
@@
The choice of weather data for the all-electric simulations has relatively little influence on grid reinforcement requirements. In the central cases presented above, we used historical weather data from 2021. An alternative approach would be to use hypothetical future weather data that incorporates the effects of climate change. Climate scientists affiliated with the Intergovernmental Panel on Climate Change have developed frameworks known as the Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP) 42. These frameworks provide both quantitative and qualitative descriptions of various societal and environmental development pathways through 2100. For example, in the middle-of-the-road scenario SSP2-RCP4.543, global mean surface temperature rises about 2 C by 2050. In a sensitivity study that used SSP2-RCP4.5 weather data for 2050, nationwide grid reinforcement requirements decrease by about 4%. While climate change increases electricity demand peaks from air conditioning in hot regions, climate change also decreases peaks from space heating in cold regions. In the sensitivity study, the latter effect outweighs the former.
@@
The findings of this paper could inform decisions about energy system design and operation. Power grid operators might use the spatially-resolved estimates of electricity demand profiles under all-electric scenarios, or the open-source software that generated them, in their capacity expansion planning. Policymakers or utility program administrators might use the demand-side management results to shape electrification incentives. For example, the perhaps surprising effectiveness of device coordination in mitigating grid impacts might warrant making electrification incentives contingent on installing equipment with the sensing, communication, and computing capabilities required for device coordination. Utility regulators might refer to the physical or economic estimates this paper when evaluating utility requests for rate increases to cover grid reinforcement costs driven by electrification. Finally, the main takeaways of this paper might influence national conversations about the future of United States energy systems.
@@
One challenge in this study was finding reliable, spatially-resolved data on distribution grid topologies and reinforcement prices. Due to a lack of topological data, we did not analyze the impacts of electrification on power quality, such as voltage support or frequency regulation. This is one possible direction for future work. Due to a lack of spatially-resolved grid reinforcement price data, we used the same price distribution in every United States county. Higher-resolution price data could enable more precise estimation of distribution grid reinforcement costs.
There are many other opportunities to extend this work. This paper’s scope covers distribution grid reinforcement driven by electrification of the residential sector and personal vehicles in the United States. Future research could extend this work to include transmission grid reinforcement and generation capacity expansion; to include electrification of the commercial and industrial sectors; or to include other countries. In particular, estimating transmission and generation costs from residential electrification would be straightforward and useful extensions. Future research could also extend the open-source EDGIE toolbox to incorporate solar photovoltaics, stationary batteries, or thermal energy storage. Finally, while this paper estimated the value of demand-side management strategies in mitigating distribution grid reinforcement costs, these strategies could also reduce transmission and generation capacity expansion costs, greenhouse gas emissions, air pollution, and household energy bills. An interesting direction for future work is to identify location-specific mixes of grid upgrades and demand-side management measures that appropriately balance the various costs and benefits.
@@
Table 1 summarizes the data sources drawn used in this paper. Distributions of building and appliance properties follow the ResStock database 44. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships 45. We generate domestic hot water time series from 31. The number of vehicles per household in each county comes from census data 46. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from 47. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to 48; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county 49. The Supplementary Information document contains further details on modeling, simulation, and data sources.
@@
Building types and floor areas NREL’s ResStock 23, 50
Business-as-usual appliances NREL’s ResStock 23, 50
Building insulation and air sealing ASHRAE code 51
Design heating and cooling temperatures EnergyStar 33
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships 45
Domestic hot water profiles Building America analysis spreadsheets 31
Commute times Census survey 52
Vehicles per household Census survey 52
Electric vehicle parameters Manufacturer specification and journals 48, 53
Weather data Oikolab 29
Climate zones ASHRAE/IECC 54
Grid upgrade prices Journals 37
@@
$0\leq q_{1}(t)\leq\eta_{1}(\theta_{1}(t))\overline{p}_{1h}+\overline{p}_{1r}.$
Here $t$ (h) denotes time, $T_{1}$ (C) is the indoor air temperature, $C_{1}$ (kWh/C) is the indoor air’s thermal capacitance, $\theta_{1}$ (C) is the outdoor air temperature, $R_{1}$ (C/kW) is the thermal resistance between indoor and outdoor air, $q_{1}$ (kW) is the thermal power supplied by the heat pump and/or resistance heater, $w_{1}$ (kW) is the exogenous thermal power from the sun, plug loads, lights, body heat, etc., $\eta_{1}$ is the heat pump’s coefficient of performance curve, $\overline{p}_{1h}$ (kW) is the heat pump’s electric power capacity, and $\overline{p}_{1r}$ (kW) is the electric power capacity of the resistance heater. In discrete time,
@@
but may saturate at an upper or lower capacity limit:
$p_{1}(k)=\begin{cases}0&\hat{q}_{1}(k)\leq 0\\ \hat{q}_{1}(k)/\eta_{1}(k)&0<\hat{q}_{1}(k)\leq\eta_{1}(k)\overline{p}_{1h}\\ (1-\eta_{1}(k))\overline{p}_{1h}+\hat{q}_{1}(k)&\eta_{1}(k)\overline{p}_{1h}<\hat{q}_{1}(k)\leq\eta_{1}(k)\overline{p}_{1h}+\overline{p}_{1r}\\ \overline{p}_{1h}+\overline{p}_{1r}&\eta_{1}(k)\overline{p}_{1h}+\overline{p}_{1r}<\hat{q}_{1}(k).\\ \end{cases}$ (4)
Experimental validation of the space heating model can be found in 28. Space cooling works similarly but without backup resistance heat.
@@
Figure 7: Distribution grid reinforcement cost distribution vs. discount rate, assuming an annual inflation rate of 2.5%. The shaded gray area is the 95% confidence region.
@@
Figure 8: Histogram of the all-electric peak aggregate demand of 1,000 Minnesota homes over 10,000 Monte Carlo simulations. The distribution is approximately Gaussian with a standard deviation of about 0.5% of the mean.
In principle, the grid reinforcement requirement $G$ (kW) is a random variable that depends on the randomly generated parameters in the building and machine models. However, $G$ is computed in each county by summing over 1,000 homes and about 2,000 vehicles. From the Central Limit Theorem, therefore, we expect $G$ to have an approximately Gaussian distribution with small variance. Fig. 8 shows the distribution of the peak aggregate demand for a 1,000-home neighborhood over 10,000 Monte Carlo simulations. The standard deviation of the peak aggregate demand is about 0.5% of the mean. Given that nationwide simulations require significant computing time and the variations of county-level peaks from one Monte Carlo simulation to the next are small, we conducted the 1,000-home simulations once for each county and treated the resulting grid reinforcement estimate as deterministic.
@@
We estimate the grid reinforcement cost using the prices $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$ from 37. Since 37 does not model the uncertainty associated with these prices, we assume they follow Gaussian distributions with standard deviations of 20% of the mean prices. As the total cost in Eq. (9) is a linear combination of $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$, the total cost is also Gaussian. We calculate its mean and variance analytically using standard formulas.

### v3

@@ RESULTS
Fig. 5 shows heat maps of the estimated reductions in grid reinforcement requirements in each county under the three demand-side management strategies considered here. The heat maps with less thermal demand from buildings (left) in Fig. 5 and with ground-source heat pumps (right) show significant reductions in grid reinforcement requirements, particularly in the coldest regions. The heat map with device coordination (middle) shows modest reductions in grid reinforcement requirements. However, these reductions are more evenly distributed across the country because, for example, coordinating electric vehicle charging can reduce peak demand anywhere, regardless of climate.
Fig. 6 shows nationwide grid reinforcement cost reductions from the three individual strategies and combinations thereof. These estimates use the mean grid reinforcement price of 960 \$/kW, which gives a nationwide cost of \$569 billion (\$180 per household per year) without demand-side management. Individually (top set of bars), the three strategies reduce costs by 19%, 28% and 43%, respectively. Perhaps surprisingly, device coordination – a strategy based mainly on software, sensing, and communication, rather than equipment upgrades – reduces costs more than switching to ground-source heat pumps, a significantly more hardware-intensive and expensive strategy. Cost savings from combined strategies (bottom set of bars) are not strictly additive. For example, the combination of less thermal demand from buildings (19% individually) and ground-source heat pumps (28%) reduces costs by 40%. Combining all three demand-side management strategies reduces nationwide grid reinforcement costs by 71%, to a total of \$165 billion (\$50 per household per year).
@@ DISCUSSION
This paper used bottom-up modeling to estimate the physical and economic impacts of electrification of all residential appliances and personal vehicles on distribution grids in each county of the lower 48 United States. The modeling tools are open-source; the cleaned input data and simulation results are free and public. This paper found that distribution grid reinforcement requirements depend mainly on space heating demand, which can increase peak aggregate demand by a factor of four or more in the coldest counties. Without demand-side management, distribution grid reinforcement could cost \$350 to \$790 billion nationwide, or \$2,800 to \$6,400 per household, but smart electrification could eliminate up to three-quarters of these costs.
@@
Figure 7: Left: Nationwide grid reinforcement costs vs. adoption rate of electric vehicles and heat pumps. Center: Nationwide costs under different modeling choices for space heating. At left and center, bar heights are costs at the mean grid reinforcement price; error bars span costs at the 2.5th to 97.5th price percentiles. Right: Nationwide costs at the mean reinforcement price vs. grid headroom today and in an all-electric future.
@@
This section discusses the sensitivity of nationwide distribution grid reinforcement requirements to several modeling assumptions, including heat pump efficiencies, whether households lower their heating temperature setpoints overnight, heat pump sizing protocols, the choice of weather data, the assumed distribution grid headroom, and the electrification adoption rate. Fig. 7 summarizes the results. The Supplementary Information document discusses the sensitivity analyses in more detail.
@@
Heat pump efficiencies significantly influence grid reinforcement requirements. In cold weather, heat pumps become less efficient and may require electric resistance backup to meet heat demand. The central cases presented above assumed that heat pumps in cold regions met the United States Department of Energy’s Cold-Climate Heat Pump Challenge specifications 38. However, in a sensitivity study where heat pumps in cold regions have efficiencies that are similar to today’s equipment, nationwide distribution grid reinforcement costs increase by 25%.
@@
Lowering space heating temperature setpoints overnight can increase grid reinforcement requirements. While lowering setpoints can save energy overnight, warming the thermal mass of a building back up in the morning can require high power and potentially trigger the use of resistance backup heat. In the central cases presented above, users did not alter setpoints overnight. However, in a sensitivity study where users lower setpoints overnight, with randomized but typical setpoint adjustment magnitudes and timings, nationwide distribution grid reinforcement costs increase by 31%.
@@
Heat pump sizing also influences grid reinforcement requirements, particularly in cold climates. In the central cases presented above, we sized heat pumps to meet the larger of the design heating load and the design cooling load, capping the size at the maximum capacity currently available on the market. If heating demand in very cold weather exceeded the heat pump’s capacity, electric resistance backup heat turned on to maintain space temperatures. An alternative approach would be to size heat pumps in colder climates to meet only the design cooling load, relying much more heavily on backup heat in cold weather. We simulated this alternative approach in a sensitivity study and found that it increases nationwide grid reinforcement requirements by 21%. In the coldest regions (climatic zones 5, 6 and 7), however, sizing heat pumps to the design cooling loads increases grid reinforcement costs by 33%.
@@
The sensitivities of grid reinforcement requirements to modeling choices on heat pump efficiency (25%), overnight temperature setpoint adjustments (31%), and heat pump sizing (21%) are more than additive. In a joint sensitivity study including all three of the alternative modeling choices discussed above, nationwide grid reinforcement costs by 86%.
@@
The choice of weather data for the all-electric simulations has relatively little influence on grid reinforcement requirements. In the central cases presented above, we used historical weather data from 2021. An alternative approach would be to use hypothetical future weather data that incorporates the effects of climate change. Climate scientists affiliated with the Intergovernmental Panel on Climate Change have developed frameworks known as the Shared Socioeconomic Pathways (SSP) and Representative Concentration Pathways (RCP) 48. These frameworks provide both quantitative and qualitative descriptions of various societal and environmental development pathways through 2100. For example, in the middle-of-the-road scenario SSP2-RCP4.549, global mean surface temperature rises about 2 C by 2050. In a sensitivity study with 2050 SSP2-RCP4.5 weather data, nationwide grid reinforcement costs decrease by 6%. While climate change increases electricity demand peaks from air conditioning in hot regions, climate change also decreases peaks from space heating in cold regions. In the sensitivity study, the latter effect outweighs the former.
Distribution grid headroom assumptions for the business-as-usual and all-electric scenarios influence grid reinforcement requirement estimates. The central cases presented above used a business-as-usual headroom range of 15% to 36% (randomized over counties) and an all-electric future headroom of 20% for all counties. We ran sensitivity studies, sweeping the headroom in the business-as-usual and all-electric scenarios from 0% to 100%. In the extreme case of 100% headroom in the business-as-usual scenario and 0% in the all-electric scenario – meaning today’s distribution grids are oversized by a factor of two and all-electric future grids have no safety margins – nationwide grid reinforcement costs \$215 billion. At the other extreme of 0% business-as-usual headroom and 100% all-electric headroom, nationwide grid reinforcement costs \$1.4 trillion.
The adoption rate of electric vehicles and heat pumps also influences grid reinforcement requirements. In the central case presented above, we estimated that 100% adoption of electric vehicles and heat pumps would cost \$569 billion nationwide. We ran sensitivity studies, varying the adoption rate from 0% to 100%. Nationwide grid reinforcement costs less than \$1 billion at 10% adoption and a modest \$40 billion at 30% adoption. Above 40% adoption, grid reinforcement requirements increase approximately linearly.
@@
The findings of this paper could inform decisions about energy system design and operation. Power grid operators might use the spatially-resolved estimates of electricity demand profiles under all-electric scenarios, or the open-source software that generated them, in their capacity expansion planning. Policymakers or utility program administrators might use the demand-side management results to shape electrification incentives. For example, the perhaps surprising effectiveness of device coordination in mitigating grid impacts might warrant making electrification incentives contingent on installing equipment with the sensing, communication, and computing capabilities required for device coordination. Utility regulators might refer to the physical or economic estimates in this paper when evaluating utility requests for rate increases to cover grid reinforcement costs driven by electrification. Finally, the main takeaways of this paper might influence national conversations about the future of United States energy systems.
@@
One challenge in this study was finding reliable, spatially-resolved data on distribution grid topologies and reinforcement prices. Due to a lack of topological data, we did not analyze the impacts of electrification on power quality, such as voltage and frequency regulation. We also modeled aggregate grid reinforcement over wide geographical areas; we did not capture the component-level details of utilities’ actual grid reinforcement processes, which span scales ranging from a single building’s service drop through substation transformers that serve thousands of buildings. These are possible directions for future work. Due to a lack of spatially-resolved grid reinforcement price data, we used the same price distribution in every United States county in this paper. Higher-resolution price data could enable more precise estimation of distribution grid reinforcement costs. Future work could also assess grid reinforcement costs at the level of utility service territories, rather than counties, to clarify how these costs might affect retail electricity prices.
There are many other opportunities to extend this work. This paper’s scope covers distribution grid reinforcement driven by electrification of the residential sector and personal vehicles in the United States. Future research could extend this work to include transmission grid reinforcement and generation capacity expansion; to include electrification of the commercial and industrial sectors; or to include other countries. In particular, estimating transmission and generation costs from residential electrification would be straightforward and useful extensions. Future research could also extend the open-source EDGIE toolbox to incorporate solar photovoltaics, stationary batteries, bidirectional electric vehicle charging, or thermal energy storage. Finally, while this paper estimated the value of demand-side management strategies in mitigating distribution grid reinforcement costs, these strategies could also reduce transmission and generation capacity expansion costs, greenhouse gas emissions, air pollution, and household energy bills. An interesting direction for future work is to identify location-specific mixes of grid upgrades and demand-side management measures that appropriately balance the various costs and benefits.
@@
Table 1 summarizes the data sources drawn used in this paper. Distributions of building and appliance properties follow the ResStock database 50. Heat pump coefficient of performance curves come from manufacturer data aggregated by the Northeast Energy Efficiency Partnerships 51. We generate domestic hot water time series from 36. The number of vehicles per household in each county comes from census data 52. We classify personal vehicles as either small (e.g., coupes and sedans) or large (e.g., pickup trucks and sport utility vehicles) and tune the distribution of small and large vehicles in each county to survey data from 53. We tune the driving efficiencies of small and large electric vehicles to the Tesla Model S and Ford F-150 lightning, respectively. For both vehicle types, driving efficiencies vary with outdoor temperatures according to 54; this variation reflects energy use for cabin heating and cooling, as well as temperature-dependent battery chemistry. Commute distances follow census data for each county 55. The Supplementary Information document contains further details on modeling, simulation, and data sources.
@@
Building types and floor areas NREL’s ResStock 25, 56
Business-as-usual appliances NREL’s ResStock 25, 56
Building insulation and air sealing ASHRAE code 57
Design heating and cooling temperatures EnergyStar 58
Heat pump coefficients of performance Northeast Energy Efficiency Partnerships 51
Domestic hot water profiles Building America analysis spreadsheets 36
Commute times Census survey 59
Vehicles per household Census survey 59
Electric vehicle parameters Manufacturer specification and journals 54, 60
Weather data Oikolab 34
Climate zones ASHRAE/IECC 61
Grid upgrade prices Journals 41
@@
$0\leq q_{1}(t)\leq\overline{q}_{1}(t)+\overline{p}_{1r}.$
Here $t$ (h) denotes time, $T_{1}$ (C) is the indoor air temperature, $C_{1}$ (kWh/C) is the indoor air’s thermal capacitance, $\theta_{1}$ (C) is the outdoor air temperature, $R_{1}$ (C/kW) is the thermal resistance between indoor and outdoor air, $q_{1}$ (kW) is the thermal power supplied by the heat pump and/or resistance heater, $w_{1}$ (kW) is the exogenous thermal power from the sun, plug loads, lights, body heat, etc., $\overline{q}_{1}$ (kW) is the heat pump’s thermal power capacity, and $\overline{p}_{1r}$ (kW) is the electric power capacity of the resistance heater. In discrete time,
@@
but may saturate at an upper or lower capacity limit. In this model, the total electric power $p_{1}$ (kW) used by the heat pump and resistance backup is
$p_{1}(k)=\begin{cases}0&\hat{q}_{1}(k)\leq 0\\ \hat{q}_{1}(k)/\eta_{1}(k)&0<\hat{q}_{1}(k)\leq\overline{q}_{1}(k)\\ \overline{q}_{1}(k)/\eta_{1}(k)+\hat{q}_{1}(k)-\overline{q}_{1}(k)&\overline{q}_{1}(k)<\hat{q}_{1}(k)\leq\overline{q}_{1}(k)+\overline{p}_{1r}\\ \overline{q}_{1}(k)/\eta_{1}(k)+\overline{p}_{1r}&\overline{q}_{1}(k)+\overline{p}_{1r}<\hat{q}_{1}(k),\end{cases}$ (4)
where $\eta_{1}$ is the heat pump’s coefficient of performance. Experimental validation of the space heating model can be found in 33. Space cooling works similarly but without resistance backup.
@@
Figure 8: Distribution grid reinforcement cost vs. discount rate, assuming an annual inflation rate of 2.5%. The shaded gray area is the 95% confidence region.
@@
Figure 9: Histogram of the all-electric peak aggregate demand of 1,000 Minnesota homes over 5,000 Monte Carlo simulations. The distribution is approximately Gaussian with a standard deviation of about 0.6% of the mean.
In principle, the grid reinforcement requirement $G$ (kW) is a random variable that depends on the randomly generated parameters in the building and machine models. However, $G$ is computed in each county by summing over 1,000 homes and about 2,000 vehicles. From the Central Limit Theorem, therefore, we expect $G$ to have an approximately Gaussian distribution with small variance. Fig. 9 shows the distribution of the peak aggregate demand for a 1,000-home neighborhood over 5,000 Monte Carlo simulations. The standard deviation of the peak aggregate demand is about 0.6% of the mean. Given that nationwide simulations require significant computing time and the variations of county-level peaks from one Monte Carlo simulation to the next are small, we conducted the 1,000-home simulations once for each county and treated the resulting grid reinforcement estimate as deterministic.
@@
We estimate the grid reinforcement cost using the prices $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$ from 41. Since 41 does not model the uncertainty associated with these prices, we assume they follow Gaussian distributions with standard deviations of 20% of the mean prices. As the total cost in Eq. (9) is a linear combination of $\pi_{\text{recurring}}$ and $\pi_{\text{capital}}$, the total cost is also Gaussian. We calculate its mean and variance analytically using standard formulas.